(2x+3)2over3=3
Can someone please help me. This problem is driving me crazy. Thank you

Answer by stanbon(48558)

(2x+3)^2 /3 = 3
Multiply both sides by 3 to get:
(2x + 3)^2 = 9
Rearrange:
(2x+3)^2 - 9 = 0
Factor:
(2x+3 - 3) (2x+3+3) = 0
2x(2x+6) = 0
2x = 0 or 2x+6 = 0
x = 0 or 2x = -3
x = 0 or x = -3/2
===================
Cheers,
Stan H.